Two angles of a triangle have the same measure and the third is 84 degrease greater than the measure of the other two. Find the measure of each angle
Answer by checkley79(3341)

X+X+82=180
2X=180-82
2X=98
X=98/2
X=49 DEGREES IS THE OTHER 2 ANGLES.